OpenStudy (anonymous):

is (2n)! the same as 2n(n!)?

OpenStudy (anonymous):

No, because, if they were, dividing them would yield 1. Here, \[\frac{(2n)!}{2n(n!)}=\frac{2n(2n-1)(2n-2)...(n+1)n!}{2n \times n!}\]\[=(2n-1)(2n-2)...(n+1)\neq1\]

OpenStudy (nowhereman):

at least for n > 1 ;-)

OpenStudy (anonymous):

Yes, I sort of assumed.

OpenStudy (anonymous):

how does the (2n)! turn into the (n+1)n! at the end there?

OpenStudy (nowhereman):

thats because he reduces n! and the first factor that remains is (n+1)n!

OpenStudy (anonymous):

does that mean (2n)! = 2n(n+1)n! ?

OpenStudy (anonymous):

thats not right at all, sorry.

OpenStudy (anonymous):

i am trying to make it look like n! = (n+1)n!, you know?

OpenStudy (anonymous):

\[(n+1)!=(n+1)n!\]

OpenStudy (nowhereman):

yes indeed \[(2n)! = 2n(2n-1)(2n-2)\cdots(2n-(n-1))n!\]

OpenStudy (anonymous):

oh i see how that happened now. thanks for spelling it out for me :)

OpenStudy (anonymous):

Fan us, then, bugaloo! ;)